# OmniDraft: A Cross-vocabulary, Online Adaptive Drafter for On-device Speculative Decoding

### v1

@@ 2 Related work
Compared to model finetuning on a fixed offline training set over multiple epochs, online adaptation focuses on continual learning or few-shot generalization to new data. Frameworks such as ProtoNet snell2017prototypical and MAML finn2017model address the few-shot learning problem. Robotics and reinforcement learning also offer insights with works like DAGGER ross2011reduction , RL^2 duan2016rl and PEARL rakelly2019efficient for continuous adaptation to new tasks. There are also works that emphasize online adaptation for LLMs as in hu2023meta ; tack2024online ; liang2024online . Lastly in speculative decoding, OSD liuonline explicitly optimizes for draft model online adaptation which is closest to ours.
@@ 4 Results
We refer to $\textnormal{SpD}_{DM}$ as the baseline speculative decoding which uses direct mapping between vocabularies. N-gram postprocessing (pp) refers to using the N-gram cache as a postprocessing technique without directly training on it as we proposed in Algorithm 1. N-gram hit refers to the average number of successful N-gram cache lookups that were accepted by the target per speculative decoding step.
@@ 4.3.2 Distillation Loss Comparisons
The different loss variants also provide different levels of performance on the test set as shown in Table 4. When trained only on the $\mathcal{L}_{\text{N-gram}}$, the training is very unstable. This could either be due to the number of n-grams being substantially smaller than the direct mapping tokens within most datasets, or the training requires additional constraints to direct towards a minima. Consequently, training on the combined loss provides better performance metrics across temperatures as well. The final $\mathcal{L}_{\text{DM}}$ KL + $\lambda\mathcal{L}_{\text{N-gram}}$ KL, 2, provides slightly better results on $\text{temperature}=1$ indicating the impact of the additional KL over the intersection vocabulary which is beneficial for the sampling process. However, we noticed tuning the scaling factor $\lambda$ becomes critical and hence we use Equation 1 for all of the experiments since it was much more stable across all tasks. We fix a $\lambda=0.2$ for all the tasks, across all experiments.
@@
Metrics $\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ NLL + $\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ KL +$\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ KL + $\mathcal{L}_{\text{N-gram}}$ KL $\text{Temperature}=0.01$ Acc Rate 0.090 0.368 0.375 0.376 Speedup 0.86x 1.51x 1.57x 1.57x $\text{Temperature}=1$ Acc Rate 0.070 0.271 0.273 0.275 Speedup 0.79x 1.23x 1.27x 1.31x

### v2

@@ 2 Related work
Compared to model finetuning on a fixed offline training set over multiple epochs, online adaptation focuses on continual learning or few-shot generalization to new data. Frameworks such as ProtoNet snell2017prototypical and MAML finn2017model address the few-shot learning problem. Robotics and reinforcement learning also offer insights with works like DAGGER ross2011reduction , RLˆ2 duan2016rl and PEARL rakelly2019efficient for continuous adaptation to new tasks. There are also works that emphasize online adaptation for LLMs as in hu2023meta ; tack2024online ; liang2024online . Lastly in speculative decoding, OSD liuonline explicitly optimizes for draft model online adaptation which is closest to ours.
@@ 4 Results
We refer to $\textnormal{SpD}_{DM}$ which is our implementation of SpD with direct mapping tokens only and is equivalent to UAG timor2025acceleratingllminferencelossless . N-gram postprocessing (pp) refers to using the N-gram cache as a postprocessing technique without directly training on it as we proposed in Algorithm 1. N-gram hit refers to the average number of successful N-gram cache lookups that were accepted by the target per speculative decoding step.
@@ 4.3 Ablations Studies
4.3.1 Scalability to Larger Target and Draft Models
We perform an ablation on scaling the target and draft model and compare the various metrics and whether it follows a similar trend. As shown in Table 3, we used a new family of LLMs, the Qwen2.5 series of models of larger sizes including 7B, 14B and 32B parameters. Datasets used are GSM8k and MBPP+HumanEval. We also performed analysis on a larger drafter of size 160M parameters. The performance on these larger models are consistent with our previous results, with larger target models resulting in 2x improvement on the GSM8K dataset. The Omnidraft framework demonstrates strong scalability as the size of the target LLM increases, as shown in the table. While the gap between the drafter and the target model grows with model size, the drafter remains the limiting factor. During training, the drafter progressively aligns with the target; however, once it reaches its alignment capacity, further increase in target size continues to yield speedup improvements, while the acceptance rate plateaus. Regarding the larger drafter, the Llama-160M parameter model was chosen after taking into consideration the increased latency of around 3x of the drafter model compared to Llama-68M. Overall, the larger drafter model does introduce a reduction in the overall speedup across all target models, however, due to its enhanced potential capability, the acceptance rate does improve across all the target models.
Table 3: Comparison of Cross-vocabulary Distillation Performance with Llama-68M and Llama-160M across three target models
Target Method GSM8K MBPP+HumanEval Acc Rate Speedup Acc Rate Speedup Qwen2.5 7B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.02x 0.10 0.94x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.401 1.66x 0.27 1.33x $\textnormal{SpD}_{DM}$ (160M) 0.18 0.70x 0.13 0.60x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.47 1.12x 0.32 0.90x Qwen2.5 14B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.17x 0.10 1.14x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.407 1.92x 0.272 1.57x $\textnormal{SpD}_{DM}$ (160M) 0.178 0.89x 0.13 0.84x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.472 1.40x 0.33 1.19x Qwen2.5 32B $\textnormal{SpD}_{DM}$ (68M) 0.153 1.30x 0.10 1.23x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.42 2.05x 0.274 1.71x $\textnormal{SpD}_{DM}$ (160M) 0.187 1.03x 0.133 0.97x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.49 1.62x 0.335 1.40x
4.3.2 N-gram Cache Memory Footprint
We also analyzed the various cache size at the end of training for each of the reported tasks as shown in Table 4. Overall n-gram cache size across tasks remain relatively small compared to the size of the draft or target model. This indicates potential feasibility for on-device setting given the small overhead.
Table 4: Summary of final N-gram cache size across tasks with Llama-68M drafter and Qwen2-7B target after online inference and distillation. Cache memory (MB) is derived from pympler.asizeof.asizeof().
GSM8K MBPP+HumanEval Alpaca XSUM Training Samples 7473 910 8000 4000 Cache Size (#n-grams) 5569 2238 20339 17013 Cache Memory (MB) 1.372 0.501 4.569 3.924
@@ 4.3.4 Distillation Loss Comparisons
[…]
@@
[…]
@@ B.5 Rank ablation for Cross-vocabulary online distillation of LoRA
B.6 Distribution Shift
B.6.1 Heterogeneous Dataset Shift
We also ablated on dataset distribution shift and data heterogeneity. We analyzed our training framework by starting the training on the GSM8K dataset and after around 1k steps, shifting to a new data distribution of MBPP+HumanEval Mix. As shown in Figure 5, initially, both speedup and acceptance rate show a steady upward trend, indicating that the model is learning effectively and becoming more efficient. At step 1000, we introduce a new dataset with a different distribution which causes a sharp drop in both acceptance rate and speedup, highlighting the impact of this shift. However, following this disruption, both metrics begin to gradually recover, demonstrating the model’s ability to adapt to the new data distribution over time. Additionally, when using our n-gram cache, the recovery seems to be much better and larger when compared to no cache. Finally, when comparing to the model that was solely trained on MBPP+HumanEval mix dataset, the model is able to slowly reach similar performance on both metrics. Overall, using our n-gram cache and despite the distribution shift, the model is showing resilience and adaptability to the new dataset.
Figure 5: Dataset Drift tracking during training by switching the dataset from GSM8K to MBPP+HumanEval at Step 1000 for Llama-60M vs Qwen-2.5-7B model
B.6.2 Target Model Shift
To further understand the impact of target model drift during training, We conducted empirical analysis on target switching between Qwen2.5 7B and 14B at arbitrary step intervals (e.g., 14B → 7B → 14B) As shown in Figure 6, our observations indicate that the drafter model is able to re-align seamlessly after each switch. Notably, switching to the 7B model yields reduced speedup due to its lower drafter latency, while maintaining alignment quality. This behavior is largely attributed to the fact that the target models belong to the same model family and as such the drafter is able to continually align during training. We also evaluated the impact of enabling the n-gram cache during target switching. Results show that the drafter recovers alignment quickly when the cache is enabled. However, in scenarios involving cross-family model swaps, a new n-gram cache is preferred due to differences in tokenization. In such cases, a simple distribution shift is insufficient to maintain alignment.
Figure 6: Model Drift tracking during training by switching the target model from 14B to 7B and back to 14B at step 300 and step 600, with Llama-68M as drafter
@@ D.3 N-gram learning
D.4 N-gram Cache Growth Rate
Figure 9: N-gram cache growth rate for Llama-68M drafter and Qwen2-7B target on GSM8K (4K subset) and MBPP+HumanEval datasets.
[…]
D.5 N-gram Cache Scalability
Table 14 indicates how speedup is impacted by different max cache size and cache eviction policies: Least Recently Used (LRU) that evicts the n-gram that hasn’t been accessed for the longest time, Least Frequently Used (LFU) that removes the n-gram with the lowest access frequency. “1/4" and “1/2" refer to using max cache size as a quarter and half of the full final cache size that we previously trained.
Among all eviction policies and max cache sizes, both evaluation metrics are better than using no cache. Specifically LRU shows diminishing returns over the growth of the cache. Meanwhile LFU plateaus at an earlier stage, with performance matching the full cache even with a quarter of the size. This could be due to LFU capturing most of the high frequency and important n-grams, compared to LRU which only captures the recent n-grams. In the context of on-device deployment, this implies with proper selection of cache size and eviction strategy, we can potentially attain robust performance given only constrained memory resources, highlighting the practical feasibility of our approach.
Table 14: Cache size scaling on MBPP+HumanEval
Cache Size Cache Type Acc Rate Speedup
No Cache - 0.219 1.29x
1/4 LRU 0.254 1.33x
LFU 0.259 1.36x
1/2 LRU 0.261 1.35x
LFU 0.263 1.36x
Full Cache - 0.267 1.36x